OpenStudy (anonymous):

Please help me write out this equation and solve: A number is divided by 4 and then decreased by five. The result is 7. What is the number?

OpenStudy (anonymous):

(x/4)-5=7

OpenStudy (anonymous):

now get x alone

OpenStudy (anonymous):

add five to both sides x/4=12

OpenStudy (anonymous):

then multiply each side by 4: x=48

OpenStudy (anonymous):

now check your answer by plugging it back into the original problem: (48/4)-5=7